THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 81 — 2 June 2006

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present the 81^{st} issue of the Magellanic Clouds Newsletter. Many thanks for your contributions and positive reactions; we hope that it may be a source of inspiration for your research.

Amongst the contributions this time, there are quite a few that deal with star clusters or with some aspect of stellar evolution or its end-products. Concerning the structure of the Magellanic Clouds, a very intriguing result is obtained by Ricardo Muñoz and collaborators, who demonstrate the presence of LMC stars as far out as 22^{∘} away from the centre of the LMC.

We would like to draw your attention to the possibility to post announcements in the newsletter. These are not limited to the advertisement of jobs or conferences. Think, for instance, of recently published books, databases or tools — we also welcome propositions of scientific ideas or problems that you would like to be addressed.

The next issue will be distributed on the 1^{st} of August; the deadline for contributions is the 31^{st} of July.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

Possible Detection of O vi from the LMC Superbubble N70

Charles W. Danforth^{1} and William P. Blair^{2}

1University of Colorado, USA 2Johns Hopkins University, USA

We present FUSE observations toward four stars in LMC superbubble N70 and compare these spectra to those of four comparison targets located in nearby field and diffuse regions. The N70 sight lines show O vi 1032 absorption that is consistently stronger than the comparison sight lines by ∼60%. We attribute the excess column density (N_{OVI} = 10^{14.03} cm^{−2}) to hot gas within N70, potentially the first detection of O vi associated with a superbubble. In a survey of 12 LMC sight lines, Howk et al. (2002a) concluded that there was no correlation between ISM morphology and N_{OVI}. We present a reanalysis of their measurements combined with our own and find a clear difference between the superbubble and field samples. The five superbubbles probed to date with FUSE show a consistently higher mean N_{OVI} than the 12 non-superbubble sight lines, though both samples show equivalent scatter from halo variability. Possible ionization mechanisms for N70 are discussed, and we conclude that the observed O vi could be the product of thermal conduction at the interface between the hot, X-ray emitting gas inside the superbubble and the cooler, photoionized material making up the shell seen prominently in Hα. We calculate the total hydrogen density n_{H} implied by our O vi measurements and find a value consistent with expectations. Finally, we discuss emission-line observations of O vi from N70.

Accepted for publication in ApJ, 646, July 20, 2006 Available from astro-ph/0603809

Red Giant stars in the Large Magellanic Cloud clusters

A. Mucciarelli^{1}, L. Origlia^{2}, F.R. Ferraro^{1}, C. Maraston^{3} and V. Testa^{4}

1Dipartimento di Astronomia, Università di Bologna, Italy 2INAF Osservatorio Astronomico di Bologna, Italy 3University of Oxford, UK 4INAF Osservatorio Astronomico di Roma, Italy

We present deep J,H,K_{s} photometry and accurate Color Magnitude Diagrams down to ∼18.5, for a sample of 13 globular clusters in the Large Magellanic Cloud. This data set combined with the previous sample of 6 clusters published by our group gives the opportunity to study the properties of giant stars in clusters with different ages (ranging from ∼80 Myr up to ∼3.5 Gyr). Quantitative estimates of star population ratios (by number and luminosity) in the Asymptotic Giant Branch, the Red Giant Branch and the He-clump, have been obtained and compared with theoretical models in the framework of probing the so-called phase transitions. The AGB contribution to the total luminosity starts to be significant at 200 Myr and reaches its maximum at ∼5-600 Myr, when the RGB Phase Transition is starting. At ∼900 Myr the full developing of an extended and well populated RGB has been completed. Both the occurrence of the AGB and RGB Phase Transitions are sharp events, lasting a few hundreds Myr only. These empirical results agree very well with the theoretical predictions of simple stellar population models based on canonical tracks and the fuel-consumption approach.

Accepted for publication in ApJ Available from astro-ph/0604139

Evolutionary Status of SNR 1987A at the Age of Eighteen

Sangwook Park^{1}, Svetozar A. Zhekov^{2,3}, David N. Burrows^{1}, Gordon P. Garmire^{1}, Judith L. Racusin^{1} and Richard McCray^{2}

1Penn State, USA 2Colorado, USA 3SRI, Bulgaria

∼18 yr after the supernova explosion, the blast wave of SNR 1987A is entering the main body of the equatorial circumstellar material, which is causing a dramatic brightening of the remnant. We recently reported the observational evidence for this event from our Chandra data (Park et al. 2005b; P05 hereafter). We present here the temporal evolution of the X-ray emitting shock parameters and the detailed description of the spectral and image analysis of SNR 1987A, on which P05 was based. While the remnant becomes brighter, the softening of the overall X-ray spectrum continues and is enhanced on around day ∼6200 (since the explosion). The two-component shock model indicates that the electron temperatures have been changing for the last ∼6 yr. The X-ray spectrum is now described by kT ∼ 0.3 keV and 2.3 keV thermal plasmas which are believed to characteristically represent the shock-heated density gradient along the boundary between the H ii region and the dense inner ring. As the blast wave sweeps through the inner circumstellar ring shining in X-rays, we expect that the shock parameters continue to change, revealing the density and abundance structure of the inner ring. Follow-up Chandra observations will thus uncover the past history of the progenitor’s stellar evolution. The origin of the relatively faint hard X-ray emission (E > 3 keV) from SNR 1987A is still unclear (thermal vs. nonthermal). Considering the continuous brightening of the hard band intensity, as well as the soft band flux, follow-up monitoring observations will also be essential to reveal the origin of the hard X-ray emission of SNR 1987A.

Accepted for publication in ApJ Available from astro-ph/0604201

Discovery of 14 radio pulsars in a survey of the Magellanic Clouds

R.N. Manchester^{1}, G.L. Fan^{2}, A.G. Lyne^{3}, V.M. Kaspi^{4} and F. Crawford^{5}

1Australia Telescope National Facility, CSIRO, Australia 2University of Hong Kong, China 3Jodrell Bank Observatory, University of Manchester, UK 4McGill University, USA 5Haverford College, USA

A systematic survey of the Large and Small Magellanic Clouds for radio pulsars using the Parkes radio telescope and the 20-cm multibeam receiver has resulted in the discovery of 14 pulsars and the redetection of five of the eight previously known spin-powered pulsars believed to lie in the Magellanic Clouds. Of the 14 new discoveries, 12 are believed to lie within Clouds, three in the Small Cloud and nine in the Large Cloud, bringing the total number of known spin-powered pulsars in the Clouds to 20. Averaged over all positions within the survey area, the survey had a limiting flux density of about 0.12 mJy. Observed dispersion measures suggest that the mean free electron density in the Magellanic Clouds is similar to that in the disk of our Galaxy. The observed radio luminosities have little or no dependence on pulsar period or characteristic age and the differential luminosity function is consistent with a power-law slope of −1 as is observed for Galactic pulsars.

Accepted for publication in ApJ Available from astro-ph/0604421

Age distribution of young clusters and field stars in the SMC

E.Chiosi^{1}, A. Vallenari^{2}, E. V. Held^{2}, L. Rizzi^{3} and A. Moretti^{1}

1Astronomy Department, Padova University, Italy 2INAF, Padova Observatory, Italy 3Institute for Astronomy, 2680 Woodlawn Drive, Hawai’i 96822-1897, USA

In this paper we discuss the cluster and field star formation in the central part of the Small Magellanic Cloud. The main goal is to study the correlation between young objects and their interstellar environment. The ages of about 164 associations and 311 clusters younger than 1 Gyr are determined using isochrone fitting. The spatial distribution of the clusters is compared with the H i maps, with the H i velocity dispersion field, with the location of the CO clouds and with the distribution of young field stars. The cluster age distribution supports the idea that clusters formed in the last 1 Gyr of the SMC history in a roughly continuous way with periods of enhancements. The two super-shells 37A and 304A detected in the H i distribution are clearly visible in the age distribution of the clusters: an enhancement in the cluster formation rate has taken place from the epoch of the shell formation. A tight correlation between young clusters and the H i intensity is found. The degree of correlation is decreasing with the age of the clusters. Clusters older than 300 Myr are located away from the H i peaks. Clusters and associations younger than 10 Myr are related to the CO clouds in the SW region of the SMC disk. A positive correlation between the location of the young clusters and the velocity dispersion field of the atomic gas is derived only for the shell 304A, suggesting that the cloud-cloud collision is probably not the most important mechanism of cluster formation. Evidence of gravitational triggered episode due to the most recent close interaction between SMC and LMC is found both in the cluster and field star distribution.

Accepted for publication in Astronomy & Astrophysics Available from astro-ph/0604166

Removing the Microlensing Blending-Parallax Degeneracy Using Source Variability

R.J. Assef^{1}, A. Gould^{1}, EROS Collaboration^{2}, MACHO Collaboration^{2} and OGLE Collaboration^{2}

1Department of Astronomy, Ohio State University, 140 W. 18th Ave., Columbus, OH 43210, USA 2Each collaboration’s author list is given in the paper.

Microlensing event MACHO 97-SMC-1 is one of the rare microlensing events for which the source is a variable star, simply because most variable stars are systematically eliminated from microlensing studies. Using observational data for this event, we show that the intrinsic variability of a microlensed star is a powerful tool to constrain the nature of the lens by breaking the degeneracy between the microlens parallax and the blended light. We also present a statistical test for discriminating the location of the lens based on the χ^{2} contours of the vector Λ, the inverse of the projected velocity. We find that while SMC self lensing is somewhat favored over halo lensing, neither location can be ruled out with good confidence.

Submitted to Astrophysical Journal Available from astro-ph/0604147

On the metallicity dependence of Wolf-Rayet winds

Jorick S. Vink^{1} and Alex de Koter^{2}

1Keele University, UK 2University of Amsterdam, The Netherlands

We have performed a pilot study of mass loss predictions for late-type Wolf-Rayet (WR) stars as a function of metal abundance, over a range between 10^{−5} < Z/Z_{⊙} < 10. We find that the winds of nitrogen-rich Wolf-Rayet stars are dominated by iron lines, with a dependence of mass loss on Z similar to that of massive OB stars. For more evolved, carbon-rich, WR stars the wind strength is also found to be dependent on the Fe abundance, so that they depend on the chemical environment of the host galaxy, but with a mass loss metallicity dependence that is less steep than for OB stars. Our finding that WR mass loss is Z-dependent is a new one, with important consequences for black hole formation and X-ray population studies in external galaxies. A further finding of our study is that the Z dependence of C-rich WR stars becomes weaker at metallicities below Z/Z_{⊙} < 1/10, and mass loss no longer declines once the metal abundance drops below Z/Z_{⊙} = 10^{−3}. This is the result of an increased importance of radiative driving by intermediate mass elements, such as carbon. In combination with rapid rotation and/or proximity to the Eddington limit — likely to be relevant for massive Population III stars — this effect may indicate a role for mass loss in the appearance and evolution of these objects, as well as a potential role for stellar winds in enriching the intergalactic medium of the early Universe.

Published in Astronomy & Astrophysics Available from astro-ph/0507352 and from http://www.astro.keele.ac.uk/∼jsv/

Implications of the metallicity dependence of Wolf-Rayet winds

John J. Eldridge^{1} and Jorick S. Vink^{2}

1Queen’s University Belfast, Northern Ireland 2Keele University, England

Aims: Recent theoretical predictions for the winds of Wolf-Rayet stars indicate that their mass-loss rates scale with the initial stellar metallicity in the local Universe.We aim to investigate how this predicted dependence affects the models of Wolf-Rayet stars and their progeny in different chemical environments. Methods: We compute models of stellar structure and evolution for Wolf-Rayet stars for different initial metallicities, and investigate how the scaling of the Wolf-Rayet mass-loss rates affects the final masses, the lifetimes of the WN and WC subtypes, and how the ratio of the two populations vary with metallicity. Results: We find significant effects of metallicity dependent mass-loss rates for Wolf-Rayet stars. For models that include the scaling of the mass-loss rate with initial metallicity, all WR stars become neutron stars rather than black holes at twice the solar metallicity; at lower Z, black holes have larger masses. We also show that our models that include the mass-loss metallicity scaling closely reproduce the observed decrease of the relative population of WC over WN stars at low metallicities.

Accepted for publication in Astronomy & Astrophysics Available from astro-ph/0603188 and from http://www.astro.keele.ac.uk/∼jsv/

Exploring Halo Substructure with Giant Stars XI: The Tidal Tails of the Carina Dwarf Spheroidal and the Discovery of Magellanic Cloud

Stars in the Carina Foreground

Ricardo R. Muñoz^{1}, Steven R. Majewski^{1}, Simone Zaggia^{2}, William E. Kunkel^{3}, Peter M. Frinchaboy^{1}, David L. Nidever^{1}, Denija Crnojevic^{2}, Richard J. Patterson^{1}, Jeffrey D. Crane^{1,4}, Kathryn V. Johnston^{5}, Sangmo Tony

Sohn^{1,6}, Rebecca Bernstein^{7} and Stephen Shectman^{4}

1Dept. of Astronomy, University of Virginia, USA 2Osservatorio Astronomico di Trieste, Italy 3Las Campanas Observatory, Chile 4Carnegie Observatories, USA 5Department of Astronomy, Wesleyan University, USA 6Korea Astronomy and Space Science Institute, South Korea 7Dept. of Astronomy, University of Michigan, USA

A new large-area Washington M, T_{2} + DDO51 filter survey of more than 10 deg^{2} around the Carina dSph galaxy reveals a spectroscopically confirmed power law radial density “break” population of Carina giant stars extending several degrees beyond the central King profile. Magellan telescope MIKE spectroscopy establishes the existence of Carina stars to at least 4.5 times its central King limiting radius, r_{lim} and primarily along Carina’s major axis. To keep these stars bound to the dSph would require a global Carina mass-to-light ratio of M/L ≥ 6, 300 (M/L)_{⊙}. The MIKE velocities, supplemented with ∼ 950 additional Carina field velocities from archived VLT+GIRAFFE spectra with r< ∼r_{lim}, demonstrate a nearly constant Carina velocity dispersion (σ_{v}) to just beyond r = r_{lim}, and both a rising σ_{v} and a velocity shear at still larger radii. Together, the observational evidence suggests that the discovered extended Carina population represents tidal debris from the dSph. Of 65 giant candidates at large angular radii from the Carina center for which MIKE spectra have been obtained 94% are associated either with Carina or a second, newly discovered diffuse, but strongly radial velocity-coherent (σ_{v}=9.8 km s^{−1}), foreground halo system. The fifteen stars in this second, retrograde velocity population have (1) a mean metallicity ∼ 1 dex higher than that of Carina, and (2) colors and magnitudes consistent with the red clump of the Large Magellanic Cloud (LMC). Additional spectroscopy of giant star candidates in fields linking Carina and the LMC show a smooth velocity gradient between the LMC and the retrograde Carina moving group. We conclude that we have found Magellanic stars almost twice as far (22^{∘}) from the LMC center than previously known.

Accepted for publication in ApJ Available from astro-ph/0605098

A multi-resolution analysis of the radio-FIR correlation in the Large Magellanic Cloud

A. Hughes^{1,2}, T. Wong^{2,3}, R. Ekers^{2}, L. Staveley-Smith^{2}, M. Filipovic^{2,4}, S. Maddison^{1}, Y. Fukui^{5} and N. Mizuno^{5}

1Centre for Supercomputing & Astrophysics, Swinburne University of Technology, Australia 2Australia Telescope National Facility, CSIRO, Australia 3School of Physics, University of New South Wales, Australia 4University of Western Sydney, Australia 5Department of Astrophysics, Nagoya University, Japan

We investigate the local correlation betwen the 1.4 GHz radio continuum and 60 μm far-infrared (FIR) emission within the Large Magellanic Cloud (LMC) on spatial scales between 0.05 and 1.5 kpc. On scales below ∼1 kpc, the radio-FIR correlation is clearly better than the correlation of the cold gas tracers with either the radio or the FIR emission. For the LMC as a whole, there is a tight correlation between the radio and FIR emission on spatial scales above ∼50 pc. By decomposing the radio emission into thermal and non-thermal components, however, we show that the scale on which the radio-FIR correlation breaks down is inversely proportional to the thermal fraction of the radio emission: regions that show a strong correlation to very small scales are the same regions where the thermal fraction of the radio emission is high. Contrary to previous studies of the local radio-FIR correlation in the LMC, we show that the slope of the relation between the radio and FIR emission is non-linear. In bright star-forming regions, the radio emission increases faster than linearly with respect to the FIR emission (power-law slope of ∼ 1.2), whereas a flatter slope of ∼ 0.6−0.9 applies more generally across the LMC. Our results are consistent with a scenario in which the UV photons and cosmic rays in the LMC have a common origin in massive star formation, but the cosmic rays are able to diffuse away from their production sites. Our results do not provide direct evidence for coupling between the magnetic field and the local gas density, but we note that synchrotron emission may not be a good tracer of the magnetic field if cosmic rays can readily escape the LMC.

Accepted for publication in MNRAS Available from astro-ph/0602380

Synthesis Imaging of Dense Molecular Gas in the N113 HII Region of the Large Magellanic Cloud

T. Wong^{1,2}, J. B. Whiteoak^{1}, J. Ott^{1}, Y.-N. Chin^{3} and M. R. Cunningham^{2}

1Australia Telescope National Facility, CSIRO 2School of Physics, University of New South Wales, Australia 3Department of Physics, Tamkang University, Taiwan

We present aperture synthesis imaging of dense molecular gas in the Large Magellanic Cloud, taken with the prototype millimetre receivers of the Australia Telescope Compact Array. Our observations of the N113 HII region reveal a condensation with a size of ∼6” (1.5 pc) FWHM, detected strongly in the 1–0 lines of HCO^{+}, HCN and HNC, and weakly in C_{2}H. Comparison of the ATCA observations with single-dish maps from the Mopra telescope indicates that the condensation is a massive clump of ∼10^{4} M_{⊙} within a larger ∼10^{5} M_{⊙} molecular cloud. The clump is centered adjacent to a compact, obscured HII region which is part of a linear structure of radio continuum sources extending across the molecular cloud. We suggest that the clump represents a possible site for triggered star formation. Examining the integrated line intensities as a function of interferometer baseline length, we find evidence for decreasing HCO^{+}/HCN and HCN/HNC ratios on longer baselines. These trends are consistent with a significant component of the HCO^{+} emission arising in a diffuse clump envelope and a lower HCN/HNC abundance ratio in dense cores.

Accepted for publication in ApJ Available from astro-ph/0604182

Far Ultraviolet Spectroscopic Explorer Survey of Magellanic Cloud Supernova Remnants

William P. Blair^{1}, Parviz Ghavamian^{1}, Ravi Sankrit^{1} and Charles W. Danforth^{2}

1Dept. of Physics and Astronomy, Johns Hopkins University, USA 2CASA, University of Colorado, Boulder, USA

We report the progress to date from an ongoing unbiased ultraviolet survey of supernova remnants in the Magellanic Clouds using the Far Ultraviolet Spectroscopic Explorer (FUSE) satellite. Earlier work with FUSE and other instruments has indicated that optical and/or X-ray characteristics of supernova remnants are not always good predictors of their brightness in the ultraviolet. This survey is obtaining spectra of a random large sample of Magellanic Cloud supernova remnants with a broad range of radio, optical, and X-ray properties. We proposed 39 objects in the Large Magellanic Cloud and 11 objects from the Small Magellanic Cloud, with a standard request of 10 ks per object using the FUSE 30′′ square aperture. To date, 39 objects have been observed in the survey (38 in the LMC and one in the SMC) and 15 have been detected, a detection rate of nearly 40%. Our survey has nearly tripled the number of UV-detected SNRs in the Magellanic Clouds (from 8 to 22). Because of the diffuse source sensitivity of FUSE, upper limits on non-detected objects are quite sensitive in many cases, dependent upon night observing fraction and whether stellar light contamination plays a role for a given object. Estimated total luminosities in O vi, based simply on scaling the flux at the observed positions to an entire object, span a broad range from considerably brighter to many times fainter than the inferred soft X-ray luminosities, indicating that O vi can be an important and largely unrecognized coolant in certain objects. We compare the optical and X-ray properties of the detected and non-detected objects but do not find a simple indicator for ultraviolet detectability. Non-detections may be due to clumpiness of the emission, high foreground extinction, slow shocks whose emission gets attenuated by the Magellanic interstellar medium, or a combination of these effects. The characteristics of individual detected supernova remnants are summarized in an Appendix.

Accepted for publication in ApJS Available from astro-ph/0605508

Massive Star Cluster Populations in Irregulars as Probable Younger

Counterparts of Old Metal-Rich Globular Cluster Populations in Spheroids

V. Kravtsov^{1,2}

1Instituto de Astronomía, Universidad Católica del Norte, Antofagasta, Chile 2Sternberg Astronomical Institute, Moscow State University, Moscow, Russia

Peak metallicities of metal–rich (MR) populations of globular clusters (GCs) belonging to spheroids of different mass fall within the somewhat conservative −0.7 ≤[Fe/H]≤ −0.3 range. Indeed, if possible age effects are taken into account, this metallicity range might become smaller. Irregulars, like the Large Magellanic Cloud (LMC), with longer timescales of their formation and lower star formation (SF) efficiency do not contain the old MRGCs with [Fe/H] > −1.0, but they are observed to form populations of young/intermediate–age massive star clusters (MSCs) with masses exceeding 10^{4}M_{⊙}. Their formation is widely believed to be accidental process fully depending on external factors. From analysis of data available on the populations and their hosts, including intermediate–age populous star clusters in the LMC, we find that their most probable mean metallicities fall within −0.7 ≤ [Fe/H] ≤ −0.3, as the peak metallicities of MRGCs do, irrespective of sings of interaction. Moreover, both the disk giant metallicity distribution function (MDF) in the LMC and the MDFs for old giants in the halos of massive spheroids exhibit significant increasing toward [Fe/H]≈ −0.5. That is in agreement with a correlation found between SF activity in galaxies and their metallicity. The formation of both the old MRGCs in spheroids and MSC populations in irregulars probably occurs approximately at the same stage of the host galaxies’ chemical evolution and is related to the essentially increased SF activity in the hosts around the same metallicity that is achieved very soon in massive spheroids, later in lower–mass spheroids, and much more later in irregulars. Changes in the interstellar dust, particularly in elemental abundances in dust grains and in the mass distribution function of the grains, may be among the factors regulating star and MSC formation activity in galaxies. Strong interaction and merger affecting the MSC formation play presumably additional role, although they can substantially intensify the internally regulated MSC formation process. Several implications of our suggestions are briefly discussed.

Accepted for publication in The Astronomical Journal, Vol. 132, No. 3 (September 2006) Available from astro-ph/0605592

On the Iron content of NGC 1978 in the LMC: a metal rich, chemically homogeneous cluster

F.R.Ferraro^{1}, A.Mucciarelli^{1}, E.Carretta^{2} and L.Origlia^{2}

1Dipartimento di Astronomia, Università di Bologna, Italy 2INAF Osservatorio Astronomico di Bologna, Italy

We present a detailed abundance analysis of giant stars in NGC 1978, a massive, intermediate-age stellar cluster in the Large Magellanic Cloud, characterized by a high ellipticity and suspected to have a metallicity spread. We analyzed 11 giants, all cluster members, by using high resolution spectra acquired with the UVES/FLAMES spectrograph at the ESO-Very Large Telescope. We find an iron content of [Fe/H]= −0.38 dex with very low σ_{[Fe/H]} = 0.07 dex dispersion, and a mean heliocentric radial velocity V_{r} = 293.1 ± 0.9 km/s and a velocity dispersion σ_{V_{r}} = 3.1 km/s, thus excluding the presence of a significant metallicity, as well as velocity, spread within the cluster.

Accepted for publication in ApJL Available from astro-ph/0605646

Physical parameters of rich LMC clusters from modeling of deep HST colour-magnitude diagrams

Leandro O. Kerber^{1} and Basílio X. Santiago^{1}

1Universidade Federal do Rio Grande do Sul, IF, CP 15051, Porto Alegre 91501-970, RS, Brazil

We present the analysis of deep colour-magnitude diagrams (CMDs) of five rich LMC clusters. The data were obtained with HST/WFPC2 in the F555W (∼V) and F814W (∼I) filters, reaching V555 ∼ 25. The sample of clusters is composed of NGC 1805 and NGC 1818, the youngest ones (τ < 100 Myr), NGC 1831 and NGC 1868, of intermediateage (400 < τ < 1000 Myr), and Hodge 14, the oldest (τ > 1200 Myr). We discuss and apply a statistical method for correcting the CMD for sampling incompleteness and field star contamination. Efficient use of the CMD data was made by means of direct comparisons of the observed to model CMDs. The CMD modeling process generates a synthetic Main Sequence (MS), where we introduce as model inputs the information about age, chemical composition, present day mass function (PDMF), fraction of unresolved binaries, distance modulus and light extinction. The photometric uncertainties were empirically determined from the data and incorporated into the model as well. Statistical techniques of CMD comparisons using 1 and 2 dimensions are presented and applied as an objective method to assess the compatibility between model and data CMDs. By modeling the CMDs from the central region we infer the metallicity (Z), the intrinsic distance modulus ((m − M)_{0}) and the reddening value (E(B − V)) for each cluster. We also determined the age for the clusters with τ > 400 Myr. By means of two-dimensional CMD comparisons we infer the following values: for NGC 1805, Z = 0.007 ± 0.003, (m − M)_{0} = 18.50 ± 0.11, E(B − V) = 0.03 ± 0.01; for NGC 1818, Z = 0.005 ± 0.002, (m − M)_{0} = 18.49 ± 0.14, E(B − V) ∼ 0.00; for NGC 1831, Z = 0.012 ± 0.002, log(τ/yr) = 8.70 ± 0.03, (m − M)_{0} = 18.70 ± 0.03, E(B − V) ∼ 0.00; for NGC 1868, Z = 0.008 ± 0.002, log(τ/yr) = 8.95 ± 0.03, (m − M)_{0} = 18.70 ± 0.03, E(B − V) ∼ 0.00; for Hodge 14, Z = 0.008 ± 0.004, log(τ/yr) = 9.23 ± 0.10, (m−M)_{0} = 18.51±0.13, E(B −V) = 0.02±0.02. Taking into account the uncertainties, these values are in accordance with the ones obtained applying the one-dimensional CMD analysis, adding reliability to these determinations.

Published in Astronomy and Astrophysics, Volume 435, Issue 1, May III 2005, pp.77-93

Mass segregation in rich LMC clusters from modelling of deep HST colour-magnitude diagrams

Leandro O. Kerber^{1} and Basílio X. Santiago^{1}

1Universidade Federal do Rio Grande do Sul, IF, CP 15051, Porto Alegre 91501-970, RS, Brazil

Aims: We used the deep colour-magnitude diagrams (CMDs) of five rich LMC clusters (NGC 1805, NGC 1818, NGC 1831, NGC 1868, and Hodge 14) observed with HST/WFPC2 to derive their present day mass function (PDMF) and its variation with position within the cluster.

Methods: The PDMF was parameterized as a power law in the available main-sequence mass range of each cluster, typically 0.9< ∼m/M_{⊙}< ∼2.5; its slope was determined at different positions spanning from the very centre out to several core radii. The CMDs in the central regions of the clusters were carefully studied earlier, resulting in accurate age, metallicity, distance modulus, and reddening values. The slope α (where Salpeter is 2.35) was determined in annuli by following two distinct methods: 1) a power law fit to the PDMF obtained from the systemic luminosity function (LF);

2) a statistical comparison between observed and model CMDs. In the second case, α is a free input parameter in the CMD modelling process where we incorporate photometric errors and the effect of binarity as a fraction of unresolved binaries (f_{bin} = 100%) with random pairing of masses from the same PDMF.

Results: In all clusters, significant mass segregation is found from the positional dependence of the PDMF slope: α< ∼1.8 for R ≤ 1.0 R_{core} and α ∼ Salpeter inside R = 2 ∼ 3 R_{core} (except for Hodge 14, where α ∼ Salpeter for R ∼ 4 R_{core}). The results are robust in the sense that they hold true for both methods used. The CMD method reveals that unresolved binaries flatten the PDMF obtained form the systemic LF, but this effect is smaller than the uncertainties in the α determination. For each cluster we estimated dynamical ages inside the core and for the entire system. In both cases we found a trend in the sense that older clusters have flatter PDMF, consistent with a dynamical mass segregation and stellar evaporation.

Published in Astronomy and Astrophysics, Volume 452, Issue 1, June II 2006, pp.155-162 Available from astro-ph/0604619

Conference Papers

The maximum mass of star clusters

M. Gieles^{1}, S.S. Larsen^{1,2}, M.R. Haas^{1}, R.A. Scheepmaker^{1} and N. Bastian^{3}

1Astronomical Institute, Utrecht University, The Netherlands 2European Southern Observatory, Garching bei München, Germany 3University College London, United Kingdom

If an universal untruncated star cluster initial mass function (CIMF) described by a power-law is assumed, the mass of the most massive star cluster in a galaxy (M_{max}) is the result of the size-of-sample (SoS) effect. This implies a dependence of M_{max} on the total number of star clusters (N). The SoS effect also implies that M_{max} within a cluster population increases with equal logarithmic intervals of age. This is because the number of clusters formed in logarithmic age intervals increases (assuming a constant cluster formation rate). This effect has been observed in the SMC and LMC. Based on the maximum pressure (P_{int}) inside molecular clouds, it has been suggested that a physical maximum mass (M_{max}[phys]) should exist. The theory predicts that M_{max}[phys] should be observable, i.e. lower than M_{max} that follows from statistical arguments, in big galaxies with a high star formation rate. We compare the SoS relations in the SMC and LMC with the ones in M51 and model the CLF for two cases: 1) M_{max} is determined by the SoS effect and 2) M_{max} =M_{max}[phys]=constant. The observed CLF of M51 and the comparison of the SoS relations with the SMC and LMC both suggest that there exists a M_{max}[phys] of 5 × 10^{5} M_{⊙} in M51. The CLF of M51 looks very similar to the one observed in the “Antennae” galaxies. A direct comparison with our model suggests that there M_{max}[phys]= 2 × 10^{6} M_{⊙}.

Oral contribution, published in Globular Clusters -Guides to Galaxies, Concepción (Chile), 6^{th} − 10^{th} March 2006 Available from astro-ph/0604578 and from www.astro.uu.nl/∼gieles

Carbon stars in local group dwarf galaxies: C and O abundances

R. Wahlin^{1}, K. Eriksson^{1}, B. Gustafsson^{1}, K. H. Hinkle^{2}, D. L. Lambert^{3}, N. Ryde^{1} and B. Westerlund^{1}

1Department of Astronomy and Space Physics, Uppsala University, Box 515, SE-751 20 Uppsala, Sweden 2National Optical Astronomy Observatory, PO Box 26732, Tucson, AZ 85726, USA 3Department of Astronomy, University of Texas, Austin, TX 78712, USA

We present abundances of carbon and oxygen as well as abundance ratios ^{12}C/^{13}C for a sample of carbon stars in the LMC, SMC, Carina, Sculptor and Fornax dwarf galaxies. The overall metallicities in these dwarf galaxies are lower than in the galactic disc. The observations cover most of the AGB and we discuss the abundance patterns in different regions along the AGB. The abundances are determined from infrared spectra obtained with the ISAAC spectrometer on VLT (R=1500) and the Phoenix Spectrometer on Gemini South (R=50000). The synthetic spectra used in the analysis were computed with MARCS model atmospheres. We find that the oxygen abundance is decreasing with decreasing overall metallicity of the system while the C/O ratio at a given evolutionary phase is increasing with decreasing oxygen abundance.

Oral contribution, published in Proceedings of the Eighth Torino Workshop on Nucleosynthesis in AGB Stars Available from astro-ph/0605244

Imprint of galaxy formation and evolution on globular cluster properties

Kenji Bekki^{1}

1University of New South Wales, Australia

We discuss the origin of physical properties of globular cluster systems (GCSs) in galaxies in terms of galaxy formation and evolution processes. Based on numerical simulations of dynamical evolution of GCSs in galaxies, we particularly discuss (1) the origin of radial density profiles of GCSs, (2) kinematics of GCSs in elliptical galaxies, (3) transformation from nucleated dwarf galaxies into GCs (e.g., omega Centauri), and (4) the origin of GCSs in the Large Magellanic Cloud (LMC).

Oral contribution, published in Globular clusters -Guides to galaxies, Concepción (Chile), 6^{th} − 10^{th} March 2006 Available from astro-ph/0605335

Review Paper

Local Group surveys for Planetary Nebulae

Laura Magrini^{1}

1INAF Osservatorio Astrofisico di Arcetri, Firenze, Italy

The Local Group (LG) represents the best environment to study in detail the PN population in a large number of morphological types of galaxies. The closeness of the LG galaxies allows to investigate the faintest side of the PN luminosity function and to detect PNe also in the less luminous galaxies, the dwarf galaxies, where a small number of them is expected.

A review of the results of the most recent imaging surveys in the LG is presented. Some applications of the surveys for PNe to the study of the star formation history of the host galaxies are analyzed. In addition, these new observational data are an invaluable resource for follow-up spectroscopy to derive the chemical properties of not only PNe, but also other important emission-line sources like H ii regions. These are fundamental tools for the discussion of the chemical evolution of the host galaxies, mapping the history of their chemical enrichment at different epochs. The latest results on this subject are presented.

Published in Proceeding of the IAU Symp. 234, Planetary Nebulae in Our Galaxy and Beyond, M. J. Barlow & R. H. Méndez, eds. Available from astro-ph/0605222